OpenStudy (anonymous):

3 candidates A,B and C contest an election.A gets atleast 40% of all the votes,B gets 20% of the number of votes that A gets and cannot get more that 80% of the number of votes that C gets.What is the maximum percentage of votes that A can get?

OpenStudy (anonymous):

new picture?

OpenStudy (anonymous):

hmm... should we be trying to minimize B and C's votes?

OpenStudy (anonymous):

was that directed at me? if so, yes :) if not.....still yes lol

OpenStudy (anonymous):

C can get nothing.

OpenStudy (anonymous):

LIke, lets make B get only 20% of A, even though B could get more.

OpenStudy (anonymous):

If 20% of A min = 80% of C then C in terms of A min is 25% of A min [Can anyone explain this]

OpenStudy (anonymous):

But C has to get some votes right? B cant have more than 80% of C. If C is 0, and B isnt, this fails.

OpenStudy (lalaly):

A+B(min)+C(min)=100 A+0.2A+0.25A=100( B is 80% of C so .8C=.2A) 1.45A=100 hence A can get 68.9% of votes...

OpenStudy (anonymous):

Beautiful answer ...I am impressed...

OpenStudy (anonymous):

lol, im just seeing it right now, very very nice solution :)

OpenStudy (anonymous):

last time you gave me a medal you broke my page o.O lol

OpenStudy (lalaly):

haha its not my fault x( sorryyy:D